OpenStudy (anonymous):

For f(x) = –2x + 7, find f(2).

OpenStudy (anonymous):

I think its 3

mathslover (mathslover):

good , you're right

OpenStudy (anonymous):

it doesn't look like you are asking for a derivative so plug in a 2 to where you see x's -2(2) + 7 = 3

mathslover (mathslover):

\[\large{f(x)=-2x+7}\] \[\large{f(2) = -2(2)+7=-4+7 =\color{blue}{3}}\]

OpenStudy (anonymous):

Thank you both

mathslover (mathslover):

You're welcome :)